OpenStudy (anonymous):

Could someone please help me how to get to a certain step when solving this kind of function. I will provide the information below:)

OpenStudy (anonymous):

How did this function f(x)=-6(x+4)(x-4)^3 become this f(x)=-6(x+4)(x^3-12x^2+48x-64) ?

OpenStudy (campbell_st):

well the cubic part has been expanded \[(x - 4)^3 = (x -4)(x-4)(x-4) = (x^2 - 8x + 16)(x - 4)\] and then you get \[f(x) = (x + 4)(x^3 -12x^2 + 48x - 64)\]

OpenStudy (anonymous):

Thanks for your help, I now understand:)